1. As of today, McCormick's market capitalization (E) is $14,237,510,000. Market value of equity (E), also known as market cap, is calculated using the following equation: market cap = share price x shares outstanding.

2. McCormick's book value of debt is $3,237,150,000. Book value of debt (D) is calculated as follows: book value of debt = last two-year average of current portion of long-term debt + last two-year average of long-term debt & capital lease obligation.

3. Cost of Equity = risk-free rate of return + beta of asset x (expected return of the market - risk-free rate of return)

Risk-free rate of return = 2.82 percent.

Beta = 0.30.

Market premium = (expected return of the market - risk-free rate of return) = 6 percent

4. Cost of debt = last fiscal year-end interest expense / book value of debt (D). McCormick's last fiscal year end interest expense is $95.7 million.

5. Use the effective tax rate of 25.705 percent.

Question2=Find the weight of debt = D / (E + D).
Answer #1

Weight of debt = D/(E+D) where

D = value of debt and E = value of equity

E = 14,237,510,000

D = 3,237,150,000

Weight of debt = 3,237,150,000/(3,237,150,000+14,237,510,000) = 18.52%

Note - Ideally, market value of debt and equity should be taken for calculating weight.